# Nonlinear optomechanical systems with quasi-periodic and chaotic dynamics

### v2

@@ 1 Introduction
The interaction between optical and mechanical modes in optomechanical systems [1] results in such effects as optomechanically induced transparency [2, 3, 4, 5, 6, 7, 8, 9], phonon lasers [10, 11, 12], mechanical resonator (oscillator) cooling [13, 14, 15, 16], photon blockade [17, 18, 19, 20, 21], normal-mode splitting [22, 23], nonclassical state preparation [24, 25, 26, 27], and dynamic stabilization of a “membrane-in-the-middle” mechanical oscillator in a optical cavity [28]. Introducing the two-laser driving, controllable enhancement of the single-photon optomechanical coupling in a prototypical Fabry-Perot cavity was realized [29]. Optomechanical systems have potential applications in quantum information processing [30]. The collapse and revival of mechanical and optical oscillations have been considered [31, 32]. The influence of nonlinearities of mechanical oscillators was studied in [33, 34]. The generation of higher order sidebands [35], optomechanical entanglement [36], mechanical quantum squeezing [37], optomechanically induced transparency [38, 39], normal-mode splitting [40], Kerr and cross-Kerr nonlinearities [41] were described.
The optomechanical interaction linear in mechanical displacement ($\sim x$) is usually considered. Since the optomechanical interaction is inherently nonlinear, higher order interactions should be taken into account. For example, quadratic interactions ($\sim x^{2}$) were usually implemented in optomechanical systems using membrane-in-the-middle configurations [6, 42]. Recently, the microwave optomechanical system in the self-oscillating regime was studied taking into account nonlinear interactions up to third order ($\sim x^{3}$) [43]. Such nonlinearities also result in cross-Kerr effects [41] and the collapse and revival of the optical field in optomechanical systems [32].
In recent years, a number of significant studies have been devoted to understanding the transition from classical nonlinear dynamics to chaotic dynamics in optomechanical systems (see, for example, [44, 45]). The obtained results lay the foundation for the application of mechanical micro- and nanoresonators in sensing, for generating random numbers and data encryption based on optomechanics, optomechanical logic and chaos computing.
@@ 2 Theory
where $\hat{a}$ ($\hat{a}^{{\dagger}}$) denotes the bosonic annihilation (creation) operators for the cavity mode with $[\hat{a},\hat{a}^{{\dagger}}]=1$, $[\hat{x},\hat{p}]=i$ (Planck’s constant $\hbar=1$), $\hat{x}$ and $\hat{p}$ are the position and momentum operators for the mechanical resonator, $\omega_{c}$ is the cavity frequency, $g_{n}$ is the optomechanical coupling strength, n = 1, 2, 3 for linear, quadratic and cubic interactions, respectively. Here $V_{d}$ is the optical driving term with the amplitude $\varepsilon$ and frequency $\omega_{d}$.
In the Heisenberg equations derived from the Hamiltonian (1) we will make the transition to classical analogies of operators and neglect noise sources [28, 46]. We assume that the decay rate $\kappa$ of the optical field is much larger than the resonant frequency of the mechanical oscillator. Then the optical field variables may be adiabatically eliminated [28] from these equations. Finally, on time scales longer than $\kappa^{-1}$, we obtain the following classical equation for the momentum $p$:
@@
where $\Delta=\omega_{c}-\omega_{d}$, and $\gamma$ is the phenomenologically introduced mechanical damping rate. Neglecting the mechanical damping, Eq. (2) can be represented in the form $dp/dt=-\partial U_{eff}/\partial x$, where
@@
So, the dynamics of the mechanical resonator is described by the effective potential $U_{eff}(x)$, in which an additional term is added to the initial harmonic potential. This term depends on the amplitude and frequency of the driving field, the coupling strengths and the cavity decay rate.
@@ 3 Results and discussion
Fig. 1 shows effects of the value and sign of the cubic optomechanical coupling strength $g_{3}$ and the amplitude $\varepsilon$ of the optical driving field on the shape of the effective mechanical potential $U_{eff}(x)$ (Eq. (3)). To express all constants in energy (frequency) units, it is necessary to make the following substitution: $g_{i}\to g_{i}(1/2m\omega_{m})^{i/2}$, where $(1/2m\omega_{m})^{1/2}\equiv x_{zpf}$ is the size of mechanical zero-point fluctuations. At the positive values of $g_{3}$ (Fig. 1a), the effective potential has one minimum, and the mechanical oscillator has one stable state of equilibrium. With a increase in $\varepsilon$, the minimum shifts towards higher values of $x$. When the cubic coupling strength $g_{3}$ becomes negative (Fig. 1c), the effective mechanical potential is transformed into a double-well one, and the mechanical oscillator can be in a metastable state. At $g_{3}=0$ and $g_{1}=0$ the potential has a symmetric two-minimum shape for arbitrary values of $\varepsilon$, but at $g_{1}\neq 0$, an asymmetry appears in the location of the minima (Fig. 1b). This asymmetry increases with increasing the linear optomechanical coupling strength and the amplitude of the optical driving field. In all three cases shown in Fig. 1 ($g_{3}$ is positive, zero and negative), in the absence of a driving field ($\varepsilon=0$) the potential is, of course, harmonic, i.e. equal to $m\omega_{m}^{2}\hat{x}^{2}/2$. Note that asymmetric double-well potentials similar to those discussed here are often encountered in condensed matter physics problems [47].
@@
When carrying out numerical calculations using Eqs. (2) and (4) to construct phase portraits, the time was maintained from 0 to 0.01 s. To construct Poincaré sections, strobing was carried out at moments of time $t_{n}=(2\pi/\Omega)n$, where integer values $n$ varied from 1 to 1800 over 0.1 s.
@@
It is interesting that the presence of all three types of the optomechanical interaction does not lead to chaotization of the oscillations of the mechanical resonator, as it happens in the presence of both linear and quadratic interactions (Fig. 4), or only quadratic interaction (Fig. 3).
@@ 4 Conclusion
We have studied model optomechanical systems with the photon-vibration interactions linear, quadratic, and cubic in mechanical displacements. It was found that under the conditions of adiabatic elimination of the optical field, the dynamics of the mechanical oscillator is determined by some effective potential. The transformation of this potential from single-well to double-well and bistable behavior of the mechanical resonator were demonstrated. The transition of the mechanical resonator from dynamic to chaotic behavior was shown when “switching on” linear, quadratic and cubic (or their combination) optomechanical interactions in mechanical displacements with a certain choice of other parameters of the system under study and initial conditions. Maintaining this set of parameters and initial conditions, we have found that the transition to the chaotic oscillation regime of the mechanical resonator is realized in the presence of both linear and quadratic interactions. In this case, the presence of only quadratic interaction is a boundary factor for the transition from dynamic to chaotic behavior. In any case, the transition to chaotic behavior is not realized without the presence of a quadratic term in mechanical displacements. The found features of nonlinear dynamics of optomechanical systems are significant for understanding physical processes in open quantum systems as well as for possible practical applications, including sensors based on mechanical micro- and nanoresonators, optomechanical logic and chaos computing.

### v3

@@ 1 Introduction
The interaction between optical and mechanical modes in optomechanical systems [1] results in such effects as optomechanically induced transparency [2, 3, 4, 5, 6, 7, 8, 9], phonon lasers [10, 11, 12, 13, 14, 15, 16, 17, 18], mechanical resonator (oscillator) cooling [19, 20, 21, 22], photon blockade [23, 25, 26, 27], normal-mode splitting [28, 29], nonclassical state preparation [30, 31, 32, 33], chaotically spiking attractors [34], and dynamical multistability induced by radiation pressure [35]. Introducing the two-laser driving, controllable enhancement of the single-photon optomechanical coupling in a prototypical Fabry-Perot cavity was realized [36]. Optomechanical systems have potential applications in quantum information processing [37]. The collapse and revival of mechanical and optical oscillations have been considered [38, 39]. The influence of nonlinearities of mechanical oscillators was studied in [40, 41]. The generation of higher order sidebands [42], optomechanical entanglement [43], mechanical quantum squeezing [44], optomechanically induced transparency [45, 46], normal-mode splitting [47], Kerr and cross-Kerr nonlinearities [48] were described.
The optomechanical interaction linear in mechanical displacement ($\sim x$) is usually considered. Since the optomechanical interaction is inherently nonlinear, higher order interactions should be taken into account. For example, quadratic interactions ($\sim x^{2}$) were usually implemented in optomechanical systems using membrane-in-the-middle configurations [6, 49]. Recently, the microwave optomechanical system in the self-oscillating regime was studied taking into account nonlinear interactions up to third order ($\sim x^{3}$) [50]. In Ref. [51], self-sustained oscillations and limit cycles were studied in the large-displacement regime when the frequency ${\omega_{c}}(x)$ of the cavity optical mode was not expanded in powers of displacement of the mechanical resonator. Such nonlinearities also result in cross-Kerr effects [52] and the collapse and revival of the optical field in optomechanical systems [39].
In recent years, a number of significant studies have been devoted to understanding the transition from classical nonlinear dynamics to chaotic dynamics in optomechanical systems (see, for example, [49, 53]). The obtained results lay the foundation for the application of mechanical micro- and nanoresonators in sensing [54, 55], for generating random numbers and data encryption based on optomechanics, optomechanical logic and chaos computing (see [53] and references therein).
@@ 2 Theory
where $\hat{a}$ ($\hat{a}^{{\dagger}}$) denotes the bosonic annihilation (creation) operators for the cavity mode with $[\hat{a},\hat{a}^{{\dagger}}]=1$, $[\hat{x},\hat{p}]=i$ (Planck’s constant $\hbar=1$), $\hat{x}$ and $\hat{p}$ are the position and momentum operators for the mechanical resonator, $\omega_{c}$ is the cavity frequency, $g_{n}$ is the optomechanical coupling strength, n = 1, 2, 3 for linear, quadratic and cubic interactions, respectively. Here $V_{d}$ is the optical driving term with the amplitude $\varepsilon$ and frequency $\omega_{d}$. The three types of coupling presented in the Hamiltonian (1) can be realized in an optomechanical system consisting of a drumhead nano-electro-mechanical resonator coupled to a microwave cavity [50].
In the quantum Langevin equations, obtained on the bases of the Heisenberg equations, which are derived from the Hamiltonian (1), we will make the transition to classical analogies of operators and neglect noise sources [56, 57]. We assume that the decay rate $\kappa$ of the optical field is much larger than the resonant frequency of the mechanical oscillator. Then the optical field variables may be adiabatically eliminated [57] from these equations. Finally, on time scales longer than $\kappa^{-1}$, we obtain the following classical equation for the momentum $p$:
@@
where $\Delta=\omega_{c}-\omega_{d}$, and $\gamma$ is the phenomenologically introduced mechanical damping rate. Note that at $\kappa\approx\omega_{m}$, the adiabaticity condition is violated. In this case the dynamical multistability in high-finesse micromechanical optical cavities has been discovered [35]. Beyond the adiabatic approximation, the effects of radiation pressure have also been studied in [58, 59, 60, 61, 62, 63, 64]. Neglecting the mechanical damping, Eq. (2) can be represented in the form $dp/dt=-\partial U_{eff}/\partial x$, where
@@
So, the dynamics of the mechanical resonator is described by the effective potential $U_{eff}(x)$, in which an additional term is added to the initial harmonic potential. This term depends on the amplitude and frequency of the driving field, the coupling strengths and the cavity decay rate. Since the presence of damping of mechanical oscillations did not qualitatively affect noticeably the transition of the optomechanical system to quasi-periodic and chaotic motion, we do not take into account in further consideration. This allows us to use the derived effective potential for the mechanical resonator in interpreting the results.
@@ 3 Results and discussion
Fig. 1 shows effects of the value and sign of the cubic optomechanical coupling strength $g_{3}$ and the amplitude $\varepsilon$ of the optical driving field on the shape of the effective mechanical potential $U_{eff}(x)$ (Eq. (3)). To express all constants in energy (frequency) units, it is necessary to make the following substitution: $g_{i}\to g_{i}(1/2m\omega_{m})^{i/2}$, where $(1/2m\omega_{m})^{1/2}\equiv x_{zpf}$ is the size of mechanical zero-point fluctuations. At the positive values of $g_{3}$ (Fig. 1a), the effective potential has one minimum, and the mechanical oscillator has one stable state of equilibrium. With a increase in $\varepsilon$, the minimum shifts towards higher values of $x$. When the cubic coupling strength $g_{3}$ becomes negative (Fig. 1c), the effective mechanical potential is transformed into a double-well one, and the mechanical oscillator can be in a metastable state. At $g_{3}=0$ and $g_{1}=0$ the potential has a symmetric two-minimum shape for arbitrary values of $\varepsilon$, but at $g_{1}\neq 0$, an asymmetry appears in the location of the minima (Fig. 1b). This asymmetry increases with increasing the linear optomechanical coupling strength and the amplitude of the optical driving field. In all three cases shown in Fig. 1 ($g_{3}$ is positive, zero and negative), in the absence of a driving field ($\varepsilon=0$) the potential is, of course, harmonic, i.e. equal to $m\omega_{m}^{2}{x}^{2}/2$. Note that asymmetric double-well potentials similar to those discussed here are often encountered in condensed matter physics problems [65].
@@
When carrying out numerical calculations using Eqs. (2) and (4) to construct phase portraits, the time was maintained from 0 to 0.01 s. The Poincaré section is the position of the particle in the phase space ($\dot{x},x$ ) at times ${t_{j}}=(2\pi/\Omega)j$, where integer values varied from 1 to 1800 over 0.1 s. In phase space, the evolution of a dynamic oscillatory system is described by a phase trajectory. The set of initial states of this system (the set of representative points in phase space) in the process of evolution corresponds to a set of phase trajectories. If the oscillatory process is established, then all the representative points of the phase plane are collected on an attractor in the form of a closed curve, which is called a limit cycle. Such a limit cycle on the Poincaré section corresponds to a point. The $m$-cycles correspond to $m$ points on the section, i.e. the oscillatory regime becomes quasi-periodic. The transition to a chaotic set of points on the Poincaré section indicates a transition to an oscillatory regime called dynamic chaos.
@@
It is interesting that the presence of all three types of the optomechanical interaction does not lead to chaotization of the oscillations of the mechanical resonator, as it happens in the presence of both linear and quadratic interactions (Fig. 4), or only quadratic interaction (Fig. 3). When adding the cubic optomechanical interaction to the linear and quadratic terms, the right global minimum of the effective potential for the mechanical oscillator becomes deeper and flatter. As a result, at a given low modulation (${\varepsilon_{M}}/\varepsilon=0.20007$), the mechanical oscillations are carried out within the limits of only this global minimum with the realized quasi-periodic motion and hints of a transition to chaotic behavior. A relatively small increase the modulation (${\varepsilon_{M}}/\varepsilon=0.5001$) takes the mechanical oscillator out of the global minimum and makes it oscillate in a two-minimum potential, leading to chaotization of its movement (see Fig. 6).
@@ 4 Conclusion
We have studied model optomechanical systems with the photon-vibration interactions linear, quadratic, and cubic in mechanical displacements. It was found that under the conditions of adiabatic elimination of the optical field, the dynamics of the mechanical oscillator is determined by some effective potential. The transformation of this potential from single-well to double-well and bistable behavior of the mechanical resonator were demonstrated. The transition of the mechanical resonator from dynamic to chaotic behavior was shown when “switching on” linear, quadratic and cubic (or their combination) optomechanical interactions in mechanical displacements with a certain choice of other parameters of the system under study and initial conditions. Maintaining this set of parameters and initial conditions, we have found that the transition to the chaotic oscillation regime of the mechanical resonator is realized in the presence of both linear and quadratic interactions. In this case, the presence of only quadratic interaction is a boundary factor for the transition from dynamic to chaotic behavior. In any case, the transition to chaotic behavior is not realized without the presence of a quadratic term in mechanical displacements. The found features of nonlinear dynamics of optomechanical systems are significant for understanding physical processes in open quantum systems as well as for possible practical applications, including sensors based on mechanical micro- and nanoresonators [54, 55], optomechanical logic and chaos computing (see [53] and references therein)